40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 797     Entry time: Tue Aug 5 10:23:00 2008
Author: steve 
Type: Update 
Category: SUS 
Subject: earthquake and venting effects 
atm 1, EQ
atm 2, vent 7 days later: venting kicks optic into place to be free,
PRM: LR magnet gets pushed in and it is stocked, side in free
Attachment 1: eq4h.jpg  100 kB  | Hide | Hide all
eq4h.jpg
Attachment 2: vent4hr.jpg  108 kB  | Hide | Hide all
vent4hr.jpg
ELOG V3.1.3-